# SDE Matching: Scalable and Simulation-Free Training of Latent Stochastic Differential Equations

### v1

@@ 1 Introduction
Differential equations are a natural choice for modeling continuous-time dynamical systems and have recently received significant interest in machine learning. Since Chen et al. (2018) proposed the adjoint sensitivity method for learning Ordinary Differential Equations (ODEs) in a memory-efficient manner, ODE-based approaches became popular in deep learning for density estimation (Grathwohl et al., 2019) and to model unevenly observed time series (Yildiz et al., 2019; Rubanova et al., 2019)].
However, ODEs describe deterministic systems and encode all uncertainty into their initial conditions. This limits the applicability of ODE-based approaches when modeling stochastic and chaotic processes. To address this limitation, Li et al. (2020) extended the adjoint sensitivity method to Latent Stochastic Differential Equations (SDEs).
Despite these advancements, training of ODE and SDE models is simulation-based, it relies on costly numerical integration of differential equations and backpropagation through the solutions. With training algorithms that are difficult to parallelize on modern hardware, Latent SDEs have resisted truly scaling.
@@
We introduce SDE Matching—a simulation-free framework for learning Latent SDE models. The key idea we adopt from matching-based approaches is direct access to latent posterior samples at any time step. This eliminates the need to integrate SDEs during training. The SDE Matching objective is estimated using the Monte Carlo method, achieving $\mathcal{O}(1)$ memory and time complexity. We summarize the complexity comparisons in Table 1.
@@
In Section 2, we provide background on the Latent SDE model. In Section 3, we discuss the connection between Latent SDEs and matching-based methods, which serve as motivation for SDE Matching. Then, in Section 4, we introduce the parameterization of the Latent SDE posterior process and formulate the SDE Matching training algorithm. In Section 5, we demonstrate that our method achieves comparable performance on both synthetic and real-world data while requiring significantly less computational cost for training. In Section 6 we discuss and contrast SDE Matching to related work. Finally, we conclude with a discussion of the limitations of our approach and potential directions for future work.
@@ 2 Background
The Latent SDE model, also known as Neural SDE, assumes that this series is generated constructively by the following stochastic process, referred to as the prior process. First, sample a latent variable $z_{0}\in\mathbb{R}^{D}$ from the initial prior distribution $p_{\theta}(z_{0})$. Next, infer the latent continuous dynamics $\mathbf{Z}={(z(t))}_{t\in[0,1]}$ by integrating the following SDE:
@@
The goal of the Latent SDE is to find a set of parameters $\theta$ that best fits a dataset of observed time series or sequence. Unfortunately, the posterior distribution of the latent variables, $p_{\theta}(\mathbf{Z}|\mathbf{X})$, is generally intractable. However, variational inference can be used to train the Latent SDE. Similar to the prior process, we introduce an approximate posterior process, which is conditioned on the observations $\mathbf{X}$. The posterior process consists of two components: the initial posterior distribution $q_{\varphi}(z_{0}|\mathbf{X})$ and a conditional SDE:
@@ 3 Diffusion Models as Latent SDEs
A key property of the denoising score matching objective is that it can be learned in a simulation-free manner. Thanks to the simplicity of the forward process, instead of integrating the SDE in Equation 8, we can directly sample from $q(z_{t}|x)$ and estimate the expectation in Equation 10 using the Monte Carlo method. This property distinguishes diffusion models from Latent SDEs. However, as we will see diffusion models are in fact a special case of Latent SDEs.
@@ 5 Experiments
Figure 2: NELBO (Equation 21) objective with respect to iteration step on 3D stochastic Lorenz attractor data. Results for Latent SDE trained with adjoint sensitivity method and SDE Matching. SDE Matching demonstrates approximatly $100$ times faster convergence in terms of iterations.
@@
To this end, we provide two sets of experiments for both synthetic data and real motion capture data. In all experiments, we use the same hyperparameters for both the conventional parameterization of Latent SDE and the parameterization described in Section 4. Additional discussions on parameterization and training are provided in Appendix B. When using the SDE Matching training procedure, we jointly optimize the parameters of the generative model $\theta$ in the prior $p_{\theta}(z_{0})$, drift $h_{\theta}(z_{t},t)$, diffusion $g_{\theta}(z_{t},t)$, and observation model $p_{\theta}(x_{t}|z_{t})$, as well as the parameters $\varphi$ of the posterior reparameterization function $F_{\varphi}$. All parameters are optimized jointly by minimizing the variational bound in Equation 21. We do not apply annealing or additional regularization techniques during training.
The experiments demonstrate that SDE Matching achieves comparable accuracy to simulation-based Latent SDE training, while significantly reducing computational complexity. Not only does SDE Matching reduce the computational cost of each training iteration, the parameterization of the posterior process also leads to faster convergence in terms of the number of iterations for further gains compared to alternatives.
@@ 5.1 Synthetic Datasets
The combined, per-iteration and convergence, runtime speed-up is over 500×500$\times$500 × compared to the adjoint sensitivity approach for this example.
@@ 5.2 Motion Capture Dataset
To validate SDE Matching on real-world data, we follow Li et al. (2020); Course & Nair (2023) and evaluate it on the motion capture dataset from Gan et al. (2015). This dataset consists of motion recordings from 35 walking subjects, represented as 50-dimensional time series with 300 observations each. The dataset is split into 16 training sequences, 3 validation sequences, and 4 test sequences, following the preprocessing method from Wang et al. (2007).
@@
SDE Matching also outperforms the other simulation-free approach, ARCTA Course & Nair (2023), which can be seen as a special case of SDE Matching. Notably, ARCTA requires drawing around 100 latent samples and evaluations of SDEs per batch, whereas SDE Matching requires only one. We attribute this result to the fact that ARCTA employs a less flexible posterior parameterization, where the distribution of latent states $q_{\varphi}(z_{t}|\mathbf{X})$ depends heavily on observations close to time $t$. This dependence makes it challenging to learn long-term dependencies and limits the ability to propagate learning signals effectively.
@@ 6 Related Work
Li et al. (2020) extended the adjoint sensitivity method from ODEs to SDEs. Unlike Neural ODEs, Latent (or Neural) SDEs are better suited for modeling inherently stochastic and chaotic processes.
Numerous studies have proposed modifications to objective functions, introduced regularized dynamics, and improved computational efficiency and numerical stability for both Neural ODEs (Kelly et al., 2020; Finlay et al., 2020; Kidger et al., 2021a) and Latent SDEs (Kidger et al., 2021b). However, most of these methods still rely on backpropagation through the numerical solutions of differential equations limiting their scalability.
Another line of work (Archambeau et al., 2007a, b) has proposed less computationally demanding techniques for inferring the Latent SDEs’s training objective. Closest to our work is Course & Nair (2023) that develops a method for training Latent SDEs in a simulation-free manner. This approach is restricted to posterior processes with Gaussian marginals and does not support state-dependent diffusion terms in the prior process. Additionally, this method assumes that the latent state in the posterior process depends only on a few temporally closest observations. As a result, for effective optimization it requires drawing approximately 100 latent samples, rather than 1, per batch during training. This method can be seen as a special case of SDE Matching with the corresponding restrictions on the parameterization of the prior process and posterior process.
More recently, Zhang et al. (2024) proposed a simulation-free technique for time-series modeling. However, this approach learns dynamics directly in the original data space rather than in a latent space. Furthermore, to model non-Markovian processes the authors condition the generative dynamics on a fixed set of past observations, making simulations more computationally expensive and limiting the model’s ability to capture long-range dependencies. This can be viewed as a special case of SDE Matching, where the latent space is constructed by concatenation of the most recent observations.
@@ 7 Limitations and Future Work
We believe that thanks to its simulation-free properties, SDE Matching has the potential to scale Latent SDE-based modeling to high-dimensional applications like audio and video generation and many other applications where they were previously infeasible. It could potentially also contribute to the development of simulation-free methods for training Latent ODE models and learning policies in reinforcement learning. However, we leave these directions for future work.
@@ 8 Conclusion
Through both synthetic and real-world experiments, we demonstrated that SDE Matching achieves performance comparable to existing adjoint sensitivity-based methods while significantly reducing computational complexity. The ability to estimate latent trajectories without solving high-dimensional SDEs makes our method particularly suitable for large-scale time-series and sequence modeling.
@@ A.4 Parameterization
Appendix B Implementation Details

### v2

@@ 1 Introduction
Differential equations are a natural choice for modeling continuous-time dynamical systems and have recently received significant interest in machine learning. Since Chen et al. (2018) introduced the adjoint sensitivity method for learning Ordinary Differential Equations (ODEs) in a memory-efficient manner, ODE-based approaches became popular in deep learning for density estimation (Grathwohl et al., 2019) and to model irregularly observed time series (Yildiz et al., 2019; Rubanova et al., 2019).
However, ODEs describe deterministic systems and encode all uncertainty into their initial conditions. This limits the applicability of ODE-based approaches when modeling stochastic and chaotic processes. To address these limitations, several works study Neural (or Latent) Stochastic Differential Equations (SDEs) (Movellan et al., 2002; Ha et al., 2018; Tzen & Raginsky, 2019; Peluchetti & Favaro, 2020; Hodgkinson et al., 2020). Latent SDEs have been shown to be more robust to data shifts (Oh et al., 2024), and applications are numerous, from neuroscience (ElGazzar & van Gerven, 2024) to video prediction (Daems et al., 2024).
To enable more memory-efficient learning, Li et al. (2020) extended the adjoint sensitivity method from ODEs to SDEs. Despite these advancements, training of ODE and SDE models is simulation-based, it relies on costly numerical integration of differential equations and backpropagation through the solutions. With training algorithms that are difficult to parallelize on modern hardware, Latent SDEs have till now resisted truly scaling.
@@
We introduce SDE Matching—a simulation-free framework for learning Latent SDE models. The key idea we adopt from matching-based approaches is direct access to samples from the latent posterior approximation at any time step. This eliminates the need to integrate SDEs during training. The SDE Matching objective is estimated using the Monte Carlo method, achieving $\mathcal{O}(1)$ memory and time complexity. We summarize the complexity comparisons in Table 1.
@@
    We establish a clear connection between diffusion models and Latent SDEs, which motivates the development of a more efficient, simulation-free training procedure.
  2. 2.
@@
  4. 4.
    We show that SDE Matching enables the application of Latent SDEs to high-dimensional problems, where training with the adjoint sensitivity method is computationally infeasible.
In Section 2, we provide background on the Latent SDE model. In Section 3, we discuss the connection between Latent SDEs and matching-based methods, which serve as motivation for SDE Matching. Then, in Section 4, we introduce the parameterization of the Latent SDE posterior process and formulate the SDE Matching training algorithm. In Section 5, we demonstrate that our method achieves comparable performance on both synthetic and real-world data while requiring significantly less computational and memory cost for training. In Section 6 we discuss and contrast SDE Matching to related work. Finally, we conclude with a discussion of the limitations of our approach and potential directions for future work.
@@ 2 Background
The Latent SDE model assumes that this series is generated constructively by the following stochastic process, referred to as the prior process. First, sample a latent variable ${z_{0}\in\mathbb{R}^{D}}$ from the initial prior distribution $p_{\theta}(z_{0})$. Next, infer the latent continuous dynamics $\mathbf{Z}={(z(t))}_{t\in[0,1]}$ by integrating the following SDE:
@@
The goal of the Latent SDE is to find a set of parameters $\theta$ that best fits a dataset of observed time series or sequences. Unfortunately, the posterior distribution of the latent variables, $p_{\theta}(\mathbf{Z}|\mathbf{X})$, is generally intractable. However, variational inference can be used to train the Latent SDE. Similar to the prior process, we introduce an approximate posterior process, which is conditioned on the observations $\mathbf{X}$. The posterior process consists of two components: the initial posterior distribution $q_{\varphi}(z_{0}|\mathbf{X})$ and a conditional SDE:
@@ 3 Diffusion Models as Latent SDEs
A key property of the denoising score matching objective is that it can be learned in a simulation-free manner. Due to the simplicity of the forward process, instead of integrating the SDE in Equation 8, we can directly sample from $q(z_{t}|x)$ and estimate the expectation in Equation 10 using the Monte Carlo method. This property distinguishes diffusion models from Latent SDEs. However, as we will see diffusion models are in fact a special case of Latent SDEs.
@@ 4.2 Posterior Process
We would like to emphasis that the correspondence between the posterior SDE in Equation 19 and its marginals $q_{\varphi}(z_{t}|\mathbf{X})$ is exact, not approximate.
@@ 4.3 Optimization
The full Bayesian treatment is also possible as a straightforward extension of the bound in Equation 21 for a given prior with a separate variational approximation for $\theta$.
We would like to emphasize that the SDE Matching objective is exactly the same variational bound as used in the adjoint sensitivity method. The key difference lies in the estimation strategy for the objective. We discuss the tightness of the variational bound in Equation 21 in Appendix B and empirically validate convergence to the true posterior for a linear SDE in Section C.1.
@@ 5 Experiments
Figure 2: NELBO (Equation 21) objective with respect to iteration step on 3D stochastic Lorenz attractor data. Results for Latent SDE trained with adjoint sensitivity method and SDE Matching. SDE Matching demonstrates approximatly $10$ times faster convergence in terms of iterations.
@@
To this end, we provide various experiments on both synthetic and real data. In all experiments, we use the same hyperparameters for both the conventional parameterization of Latent SDE and the parameterization described in Section 4. Additional details on parameterization and training is provided in Appendix D. When using the SDE Matching training procedure, we jointly optimize the parameters of the generative model $\theta$ in the prior $p_{\theta}(z_{0})$, drift $h_{\theta}(z_{t},t)$, diffusion $g_{\theta}(z_{t},t)$, and observation model $p_{\theta}(x_{t}|z_{t})$, as well as the parameters $\varphi$ of the posterior reparameterization function $F_{\varphi}$. All parameters are optimized jointly by minimizing the variational bound in Equation 21. We do not apply annealing or additional regularization techniques during training.
The experiments demonstrate that SDE Matching achieves comparable or better accuracy than simulation-based Latent SDE training, while significantly reducing computational complexity. Not only does SDE Matching reduce the computational cost of each training iteration, the parameterization of the posterior process also leads to faster convergence in terms of the number of iterations for further gains compared to alternatives. The findings suggest that SDE Matching enhances the scalability of Latent SDEs to longer and higher-dimensional time series.
@@ 5.1 Synthetic Datasets
The combined, per-iteration and convergence, runtime speed-up is over 50×50$\times$50 × compared to the adjoint sensitivity approach for this experiment.
To evaluate robustness with respect to time horison, we compare the gradient norms of the objective function with respect to model parameters for both SDE Matching and the adjoint sensitivity method. When integrating over time horizons $T\in{1,2,5,10}$, the adjoint sensitivity method yields the following $\log_{10}$-based gradient norms: $6.26\pm 0.14$, $6.95\pm 0.20$, $7.91\pm 0.23$ and $8.82\pm 0.28$. These results demonstrate that the gradient norms grow exponentially with the time horizon for adjoint sensitivity methods.
In contrast, SDE Matching maintains a stable gradient norm of $4.92\pm 0.24$ (in $\log_{10}$ scale) across all time horizons, indicating significantly greater stability for long time series modeling.
In Section C.2 we provide additional experiments demonstrating robustness of the SDE Matching training procedure.
@@ 5.2 Motion Capture Dataset
To validate SDE Matching on real-world data, we follow Li et al. (2020); Course & Nair (2023) and evaluate it on the motion capture dataset from Gan et al. (2015). This dataset consists of motion recordings from subject 35 walking, represented as 50-dimensional time series with 300 observations each. The dataset is split into 16 training sequences, 3 validation sequences, and 4 test sequences, following the preprocessing method from Wang et al. (2007).
@@
(a) Adjoint sensitivity method
(b) SDE Matching
Figure 3: Qualitative comparison of generated video sequences depicting a moving pendulum. Samples generated by the model trained with adjoint sensitivity method and SDE Matching with same number of training iterations.
SDE Matching also outperforms the other simulation-free approach, ARCTA Course & Nair (2023), which can be seen as a special case of SDE Matching. Notably, ARCTA requires drawing around 100 latent samples and evaluations of SDEs per batch, whereas SDE Matching requires only one. We attribute this result to the fact that ARCTA employs a less flexible prior and posterior parameterization, where the volatility function $g_{\theta}$ is state-independent and distribution of latent states $q_{\varphi}(z_{t}|\mathbf{X})$ depends heavily on observations close to time $t$. This dependence makes it challenging to learn long-term dependencies and limits the ability to propagate learning signals effectively.
An extended discussion on the importance of state-dependent volatility functions $g_{\theta}$, along with experiments on the state-dependent stochastic Lotka-Volterra system, is provided in Section C.3.
5.3 Video Dataset
Finally, we evaluate the performance of SDE Matching in a high-dimensional setting by modeling sequences of $32\times 32$ images—i.e. video data—depicting a moving pendulum.
Using an identical number of training iterations ($20$k), SDE Matching successfully learns the underlying dynamics in approximately 20 minutes on a single GPU. In contrast, the adjoint sensitivity method requires around 30 hours for the same number of iterations and still fails to accurately capture the system’s dynamics.
Visualizations of the generated sequences are provided in Figure 3.
@@ 6 Related Work
Traditionally, Monte Carlo methods have been used to train SDEs (Movellan et al., 2002; Beskos et al., 2006; van der Meulen & Schauer, 2017). Another line of work (Archambeau et al., 2007a, b) has focused on variational inference techniques for inferring the Latent SDEs’s training objective. Discretization is used by Ryder et al. (2018) to enable more expressive variational approximations, and Ha et al. (2018) draw the connection to stochastic optimal control for more flexibility.
More recently, Li et al. (2020) extended the adjoint sensitivity method from ODEs to SDEs. Unlike Neural ODEs, Latent SDEs are better suited for modeling inherently stochastic and chaotic processes and are more robust to data shifts.
Many studies have proposed modifications to objective functions, introduced regularized dynamics, and improved computational efficiency and numerical stability for both Neural ODEs (Kelly et al., 2020; Finlay et al., 2020; Kidger et al., 2021a) and Latent SDEs (Kidger et al., 2021b). However, most of these methods still rely on backpropagation through the numerical solutions of differential equations limiting their scalability. In Verma et al. (2024) the authors instead rely on a site-based Gaussian approximation and suggest a fixed-point training algorithm.
Closest to our work is Course & Nair (2023) that develops a method for training Latent SDEs in a simulation-free manner. This approach is restricted to posterior processes with Gaussian marginals and does not support state-dependent diffusion terms in the prior process. Additionally, this method assumes that the latent state in the posterior process depends only on a few temporally closest observations. As a result, for effective optimization it requires drawing approximately 100 latent samples, rather than 1, per batch during training. This method can be seen as a special case of SDE Matching with the corresponding restrictions on the parameterization of the prior process and posterior process.
Recently, Zhang et al. (2024) proposed a simulation-free technique for time-series modeling. However, this approach learns dynamics directly in the original data space rather than in a latent space. Furthermore, to model non-Markovian processes the authors condition the generative dynamics on a fixed set of past observations, making simulations more computationally expensive and limiting the model’s ability to capture long-range dependencies. This can be viewed as a special case of SDE Matching, where the latent space is constructed by concatenation of the most recent observations.
@@ 7 Limitations and Future Work
We believe that thanks to its simulation-free properties, SDE Matching has the potential to scale Latent SDE-based modeling to high-dimensional applications like audio and video generation and many other applications where they were previously infeasible. It could potentially also contribute to the development of simulation-free methods for training Latent ODE models and learning policies in reinforcement learning. We leave these directions for future work.
@@ 8 Conclusion
Through both synthetic and real-world experiments, we demonstrated that SDE Matching achieves performance comparable to or better than existing adjoint sensitivity-based methods, while significantly reducing computational complexity. The ability to estimate latent trajectories without solving high-dimensional SDEs makes our method particularly suitable for large-scale time-series and sequence modeling.
@@ A.4 Parameterization
Appendix B Theoretical Considerations
In this section, we provide an extended discussion on the connection between the prior and posterior processes. Our primary goal is to address the question: under what conditions does the variational bound in Equation 21 become tight and match the true likelihood of the model?
The variational bound becomes tight when the approximate posterior distribution exactly matches the true posterior. In the case of the Latent SDE, this means that the posterior process must match the true posterior of the prior process.
The prior process (Equation 15) is given by:
$dz_{t}$ $=h_{\theta}(z_{t},t)dt+g_{\theta}(z_{t},t)d\mathbf{w}.$ (39)
Given a series of observations $\mathbf{X}$, the true posterior process can be derived via Doob’s $h$-transform (Mider et al., 2021; Park et al., 2024), yielding:
$dz_{t}$ $=\Big{[}\underbrace{h_{\theta}(z_{t},t)+g_{\theta}(z_{t},t)g^{\top}_{\theta}(z_{t},t)\nabla_{z_{t}}\log p_{\theta}\left(\mathbf{X}_{t}|z_{t}\right)}_{h_{\theta}(z_{t},t,\mathbf{X})}\Big{]}dt+g_{\theta}(z_{t},t)d\mathbf{w},$ (40)
where $\mathbf{X}_{t}=\{x_{s}:x_{s}\in\mathbf{X},s\geq t\}$ denotes the set of future observations from time $t$ onward.
We can also consider the corresponding deterministic process that follows the posterior marginals $p_{\theta}(z_{t}|\mathbf{X})$
$\frac{dz_{t}}{dt}$ $=\bar{h}_{\theta}(z_{t},t,\mathbf{X})$ (41)
$=h_{\theta}(z_{t},t)+g_{\theta}(z_{t},t)g^{\top}_{\theta}(z_{t},t)\left[\nabla_{z_{t}}\log p_{\theta}\left(\mathbf{X}_{t}|z_{t}\right)-\frac{1}{2}\nabla_{z_{t}}\log p_{\theta}\left(z_{t}|\mathbf{X}\right)\right]-\frac{1}{2}\nabla_{z_{t}}\cdot\left[g_{\theta}(z_{t},t)g^{\top}_{\theta}(z_{t},t)\right]$ (42)
Now, consider the approximate posterior process (Equation 19) of the form:
$dz_{t}$ $=f_{\theta,\varphi}(z_{t},t,\mathbf{X})dt+g_{\theta}(z_{t},t)d\mathbf{w}.$ (43)
The approximate process matches the true posterior process if the drift terms are equal: $f_{\theta,\varphi}(z_{t},t,\mathbf{X})\equiv h_{\theta}(z_{t},t,\mathbf{X})$ or equivalently, if the corresponding deterministic processes (Equations 17 and 42) match: $\bar{f}_{\theta,\varphi}(z_{t},t,\mathbf{X})\equiv\bar{h}_{\theta}(z_{t},t,\mathbf{X})$. In terms of the reparameterization function $F_{\varphi}(\varepsilon,t,\mathbf{X})$ (Equation 16) that defines the approximate posterior process, tthe variational bound becomes tight, and the approximate posterior process matches the true posterior, if $F_{\varphi}$ learns the solution trajectories of the ODE defined by $\bar{h}_{\theta}(z_{t},t,\mathbf{X})$ (Equation 42).
Unfortunately, as is typical in variational inference, the true posterior is intractable. Nevertheless, its analytical form provides valuable insight into the behavior that the approximate posterior should approximate. Therefore, with a sufficiently flexible reparameterization function $F_{\varphi}(\varepsilon,t,\mathbf{X})$, the SDE Matching approach should, in principle, be capable of learning the true posterior process and the variational bound becomes equal to the negative log-marginal likelihood.
Appendix C Additional Results
C.1 Convergence to Likelihood
To empirically validate the tightness of the ELBO in SDE Matching, as discussed in Appendix B, we set up a linear stochastic system of the form:
$dz_{t}$ $=F(t)z_{t}dt+L(t)d\mathbf{w}$ (44)
$x_{t}$ $=H(t)z_{t}+r_{t},$ (45)
where $r_{t}\sim\mathcal{N}(0,R(t))$. For this system, the true log-marginal likelihood can be computed using the Kalman filter (Sarkka & Solin, 2019).
In this experiment, we use simple coefficients: $F(t)=-t$, $L(t)=t$, $H(t)=1$, and $R(t)=0.01$. The model is parameterized identically to our experiment on the 3D stochastic Lorenz attractor (Section 5.1). As shown in Figure 4(a), the SDE Matching ELBO closely converges to the true log-marginal likelihood. Furthermore, similar to our results in Section 5.1, SDE Matching achieves significantly faster convergence than the adjoint sensitivity method and requires approximately 20×20$\times$20 × less time per training iteration.
C.2 Robustness of Gradients
We present additional experiments to assess the robustness of SDE Matching.
Using the setup from Section C.1, we compare the gradient norms and variances over training time. As shown in Figure 4, SDE Matching consistently exhibits both lower gradient norms and reduced variance compared to the adjoint sensitivity method.
C.3 Importance of State-Dependent Diffusion Term
We also present an example that highlights the importance of state-dependent diffusion term (volatility), a feature not supported by the previous simulation-free approach ARCTA (Course & Nair, 2023).
To demonstrate this, we design an experiment based on the stochastic Lotka–Volterra system with dynamics:
$dx$ $=(\alpha x+\beta xy)dt+\sigma xd\mathbf{w}$ (46)
$dy$ $=(\delta xy+\gamma y)dt+\sigma yd\mathbf{w}.$ (47)
We follow the setup from Course & Nair (2023) (Appendix G.1) with volatility coefficient $\sigma=0.15$ and use the same parameterization as in Section 5.1.
We then apply SDE Matching to train models with both state-independent and state-dependent volatility functions. As illustrated in Figure 5, the model with state-independent volatility fails to accurately capture the correct trajectory structure, whereas the state-dependent model succeeds.
State-dependant volatility is often a relevant property of many dynamical systems in e.g. finance (Oksendal, 2003), biology (Vadillo, 2019), neuroscience (ElGazzar & van Gerven, 2024), and causal inference (Peters et al., 2022), where Latent SDEs can be applied for modelling. The Lotka–Volterra system which we use in this experiment is an example of such a biological system. In Neural SDEs, state-dependant volatility also leads to dynamics that is more robust to distribution shift (Oh et al., 2024). Finally, the state-dependant volatility is an important contributing factor in our MOCAP experiment (Section 5.2), where the state-dependent SDE Matching and Li et al. (2020) both perform much better than the state-independent ARCTA (Course & Nair, 2023). Therefore, we believe that, together with more general marginals of the posterior process through normalising flows, enabling simulation-free training with state-dependant volatility is also an important contribution of SDE Matching.
(a) Convergence of NELBO
(b) Norm of gradient
(c) Variance of gradients
Figure 4: Linear system training dynamics for Latent SDE trained with adjoint sensitivity method and SDE Matching.
Figure 5: Lotka–Volterra system, learned trajectories for SDE Matching with state-independent and state-dependent volatility.
Appendix D Implementation Details